The perimeter of a rectangle is 210m. If the length of this rectangle is three-fourths of its width, what are the dimensions?
Answer by checkley79(3341)

L=3/4W
2L+2W=210
2(3W/4)+2W=210
6W/4+2W=210
(6W+4*2W)/4=210
(6W+8W)/4=210
14W/4=210 CROSS MULTIPLY
14W=210*4
14W=840
W=840/14
W=60 ANS. FOR THE WIDTH.
L=(3/4)60
L=180/4
L=45 ANS. FOR THE LENGTH.
PROOF:
2*45+2*60=210
90+120=210
210=210